OpenStudy (anonymous):

(10)/x(x-2) + (4)/(x) = (5)/(x-2)

OpenStudy (anonymous):

Multiply throughout to get rid of the fractions then solve: 10 / x(x - 2) + 4 / x = 5 / (x - 2) 10 + 4(x - 2) = 5x 10 + 4x - 8 = 5x x = 2